OpenStudy (anonymous):

5^(x+2)-25^(x+0.5)=2(5^(x+1)), solve the equation leaving answer in 2 decimal places. I used substitution and arrive at the answer of a=0 & a=3, but do not know how to get x.

OpenStudy (dumbcow):

where did the a come from?

OpenStudy (dumbcow):

i get x = 0.68

OpenStudy (dumbcow):

By rearranging and expanding terms: \[\rightarrow \frac{25*5^{x} -5*5^{x}*5^{x}}{5*5^{x}} = 2\] After factoring out a 5*5^x \[\rightarrow 5-5^{x} = 2\] \[5^{x} = 3\]